OpenStudy (anonymous):

Which number replaces the box to make this number sentence true? 7 × (5 – 2) = (7 × 5) – ( × 2)

OpenStudy (anonymous):

Do yuo have any idea how to solve this ?

OpenStudy (anonymous):

*you

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok. Do you know how to do distribution property?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So it says 7*(5-2) That is the same as 7(5-2) Do you know how much that equals

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(7*5)-(7*2) Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now look at the question It says (7*5)-(__*2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Give me a second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok Yes so what do you think will go in the blank by looking at the answer i gave you and the question i just gave you

OpenStudy (anonymous):

7

OpenStudy (anonymous):

You are Absolutely _________

OpenStudy (anonymous):

CORRECT!! :) :) :) :) :) :) :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so the answer is 7*(5-2)=(7*5)-(7*2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Anything else

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Which number replaces the box to make this number sentence true? 7 × (3 + 9) = 21 +

OpenStudy (anonymous):

Tell me in another thread and close this question. (Do you know how to medal cause it looks like you are new?)

OpenStudy (anonymous):

im new

OpenStudy (anonymous):

Do you now how to medal?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

So you see Best Response on the right side of every answer? (not yours because you cant say yours is the best) If you think that my answer was the best and you really like how i solve or help you then you give me a medal by clicking on Best Response

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

your well come

OpenStudy (anonymous):

Which number replaces the box to make this number sentence true? 7 × (3 + 9) = 21 +

OpenStudy (anonymous):

63

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

gotta go bu if you have questions just message me.

OpenStudy (anonymous):

ok bye

OpenStudy (anonymous):

ok bye